OpenStudy (anonymous):

How do I verify the identity of this? tan(x) + cot(x) = 1 / (sin(x) * cos(x)) I know each side equals csc(x) * sec(x) but how do I get the answer? @jim_thompson5910 Could you please help me?

OpenStudy (anonymous):

All I know is that I can make the left side equal to (sin(x) / cos(x)) + (cos(x) / sin(x))

OpenStudy (whpalmer4):

Okay, why don't you make a common denominator on the left side?

OpenStudy (anonymous):

Both sides need to have a denominator of sin(x) * cos(x)?

OpenStudy (whpalmer4):

One thing I like to do when doing trig identities is to make a substitution like c = cos (x) s = sin (x) just makes the writing a bit easier, then when I'm done with the algebra (like adding these two fractions) I undo the substitution. same with rationalizing radicals...

OpenStudy (anonymous):

I meant both parts of the left.

OpenStudy (whpalmer4):

So your problem would look like \[\frac{s}{c} + \frac{c}{s} = \frac{1}{c*s}\]

OpenStudy (whpalmer4):

If nothing else, it's a lot easier to type into OS :-)

OpenStudy (whpalmer4):

So to make a common denominator on the left side, what would you do?

OpenStudy (anonymous):

(sin(x) / cos(x)) + (cos(x) / sin(x)) would become (cos(x) * sin(x) / cos(x) * sin(x)) + (cos(x) * sin(x) / sin(x) * cos(x)) Sorry, I'm very rusty on my fraction skills.

OpenStudy (anonymous):

You have \[\sin/\cos + \cos/\sin\] just get a common denominator and simplify the numerator. You'll see it when you get it in one fraction :)

OpenStudy (whpalmer4):

Let's just do it with the letters, like I suggested: \[\frac{s}{c} + \frac{c}{s} = \frac{1}{c*s}\]Multiply the top and bottom of the left fraction by the denominator of the right fraction: \[\frac{s}{c}*\frac{s}{s} + \frac{c}{s} = \frac{1}{c * s}\]Now do the same for the right fraction: \[\frac{s}{c}*\frac{s}{s} + \frac{c}{s}*\frac{c}{c}= \frac{1}{c * s}\]Combine the two fractions \[\frac{s^2+c^2}{c*s} = \frac{1}{c * s}\]Now we can just eliminate the denominator \[s^2+c^2=1\]And if we substitute back our trig functions, it becomes\[\sin^2 x+ \cos^2 x = 1\]And hopefully you recognize that very basic identity!

OpenStudy (anonymous):

Thanks! I understand it now. I never thought it was that simple ha ha. Do you mind helping me with another question?